OpenStudy (anonymous):

the no. of values of K for which the system of equations (K-1)x+(3K+1)y+2Kz = 0, (K-1)x+(4K-2)y+(K+3)z=0 and 2x+(3k+1)y+(3K-1)z=0 has a common non zero solution is.. (a) 0 (b) 1 (c) 2 (d) 3

OpenStudy (apoorvk):

Okay, in here you gotta use Cramer's rule. (studied that yet? (matrices, determinants)

OpenStudy (earthcitizen):

(k-1)x+(3k+1)y+2kz=0 (k-1)x+(4k-2)y+(k+3)z=0 2x+(3k+1)y+(3k-1)z=0 \[\left[\begin{matrix}(k-1) & (3k+1)&2k \\ (k-1) & (4k-2)& (k+3)\\2x &(3k+1)&(3k-1)\end{matrix}\right]\left(\begin{matrix}x \\ y\\z\end{matrix}\right)=\left(\begin{matrix}0 \\ 0\\0\end{matrix}\right)\] let Au=b, where A=\[\left[\begin{matrix}(k-1) & (3k+1)&2k \\ (k-1) & (4k-2)& (k+3)\\2x &(3k+1)&(3k-1)\end{matrix}\right]\] u=\[\left(\begin{matrix}x \\ y\\z\end{matrix}\right)\] b=\[\left(\begin{matrix}0 \\ 0\\0\end{matrix}\right)\] To find u. Use,\[u=A^-1b\] recall, \[A^-1 = (1/detA)(adjA)\] then solve

OpenStudy (earthcitizen):

Gaussian elimination can be used here

OpenStudy (earthcitizen):

\[ \det A = 6 k^3-34 k^2+46 k+6 = 0\] \[k = 3, (4 - \sqrt19)/3, (4 + \sqrt19)/3\]

OpenStudy (earthcitizen):

Hence, k =3 we can find x,y and z \[\left[\begin{matrix}2 & 10&6 \\ 2 & 10&6\\2&10&8\end{matrix}\right]\left(\begin{matrix}x \\ y\\z\end{matrix}\right)=0\] x=0, y=0, z=0

OpenStudy (anonymous):

thx. :)

OpenStudy (earthcitizen):

is this correct ?

OpenStudy (earthcitizen):

then again i used Gaussian elimination and reduced row echelon augumented matrix. k = 1, x=0, y=0 and z=0